A constant force is applied to an object, causing the object to accelerate at 5.50 m/s2 . What will the acceleration be if

The force is doubled?

The object's mass is doubled?

The force and the object's mass are both doubled?

The force is doubled and the object's mass is halved?
Answer #1

Given that

force F1=F

acceleration a1=5.5 m/s^2

now we find the acceleration if force is doubled

F1/F2=a1/a2

F/2F=5.5/a2

acceleration a2=11 m/s^2

now we find the acceleration if mass is doubled

m1a1=m2a2

m*5.5=2m*a2

acceleration a2=2.75 m/s^2

now we find the find the acceleration if force and mass is doubled

F1/F2=m1a1/m2a2

F/2F=m*5.5/2m*a2

acceleration is not change

now we find the acceleration if force is doubled and mass is halved

F1/F2=m1a1/m2a2

F/2F=m*5.5/(m/2)a2

acceleration a2=22 m/s^2